Hello, I need help with this problem:
The intensity of a light source varies inversely with the square of the distance from the source. Four feet from the source the intensity is 9 foot candles. What is the intensity 3 feet from the source.
Thank you.

Answer by stanbon(75887)

The intensity of a light source varies inversely with the square of the distance from the source.
---
I = k/d^2
---
Solve for "k" using "Four feet from the source the intensity is 9 foot candles".
---
9 = k/4^2
k = 9*16 = 144
----
Equation:
I = 144/d^2
-----
What is the intensity 3 feet from the source.
---
I = 144/3^2
I = 16 ft. candles
======================
Cheers,
Stan H.